### 19 Mar 2024 v1

# Multimodal Fusion Method with Spatiotemporal Sequences and Relationship Learning for Valence-Arousal Estimation

Abstract
This paper presents our approach for the VA (Valence-Arousal) estimation task in the ABAW6 competition. We devised a comprehensive model by preprocessing video frames and audio segments to extract visual and audio features. Through the utilization of Temporal Convolutional Network (TCN) modules, we effectively captured the temporal and spatial correlations between these features. Subsequently, we employed a Transformer encoder structure to learn long-range dependencies, thereby enhancing the model’s performance and generalization ability. Our method leverages a multimodal data fusion approach, integrating pre-trained audio and video backbones for feature extraction, followed by TCN-based spatiotemporal encoding and Transformer-based temporal information capture. Experimental results demonstrate the effectiveness of our approach, achieving competitive performance in VA estimation on the AffWild2 dataset.
1 Introduction
With the continuous development of artificial intelligence technology and society, human-computer interaction has become a research field of great concern. An excellent human-computer interaction system not only needs to have efficient functionality but also needs to consider the user’s emotions and experiences. Emotion analysis plays an important role in this regard, as it is expected to provide a more accurate understanding of human emotions, thereby designing human-computer interaction systems that are more user-friendly and closer to user needs.
In recent years, human emotional behavior analysis has received increasing attention. Commonly used human expression modalities in this field include Action Units (AU), basic expression categories (EXPR), and Valence-Arousal (VA). Action Units describe the movement of specific facial regions and are the smallest units for describing expressions. Basic expression categories classify expressions into several emotional categories such as happiness, sadness, etc. VA is a model that includes two continuous values representing the valence and arousal of emotions, which can better describe human emotional states.
The ABAW workshop aims to address the challenges of in-the-wild emotional behavior analysis, which reflects a key characteristic of real-world Human-Computer Interaction (HCI) systems. The objective is to develop intelligent machines and robots capable of understanding human emotions, moods, and behaviors. By interacting with humans in a ”human-centered” manner, these systems provide engaging experiences and effectively serve as digital assistants. Such interactions should not be constrained by individual backgrounds, age, gender, race, education level, occupation, or social status. Therefore, the development of intelligent systems capable of accurately analyzing human behavior in the wild is essential for fostering trust, understanding, and intimacy between humans and machines in real-life environments. The advancement of this technology holds promise for the future of HCI, enabling more sophisticated and human-centric experiences in everyday life. To promote the development of human emotional behavior analysis, the Affective Behavior Analysis in-the-wild (ABAW) competition has proposed a series of challenges aimed at overcoming obstacles in this field. The ABAW competition has constructed large-scale multimodal video datasets, such as Affwild [29, 14]and Affwild2[12, 13], which provide valuable resources for researchers, greatly advancing the progress of in-the-wild facial expression analysis and accelerating the application in related industries. The Affwild2 dataset contains a large number of videos, most of which are annotated frame by frame with AU, basic expression categories, VA, and other labels, providing researchers with rich annotated data.
Emotion recognition[26, 10], as a critical task, plays a crucial role in human emotion understanding and human-computer interaction. Among them, the estimation of Valence and Arousal is a key component of emotion recognition. Through the dimensional model[23], we can view emotional states as points in continuous space, with Valence and Arousal serving as axes. However, achieving accurate VA estimation faces numerous challenges. Firstly, the expression of emotions is subjective and varies between individuals, leading to uncertainty in emotion recognition results due to differences in how different individuals perceive emotions. Secondly, emotional expression is typically dynamic and influenced by time, necessitating consideration of long-term temporal dependencies to better capture the process of emotional change. Additionally, emotional expression takes various forms and may be influenced by external factors and individual experiences, requiring emotion recognition models to have a certain level of robustness to adapt to different contexts and individual characteristics. Addressing these challenges requires integrating multimodal data and advanced models to improve the accuracy and robustness of VA estimation, thereby advancing the application and development of emotion recognition technology in the fields of human-computer interaction and affective computing.
To address the challenges above, we employ a multimodal data fusion approach. Firstly, we utilize a pre-trained audio VGGnet as the backbone for extracting dynamic audio features from VGGish[9], and an IResnet[6] pre-trained with the arcface[5] method as the backbone for extracting video frames. Consequently, the expressiveness of deep features in both visual and audio modalities can be further enhanced through fine-tuning. Subsequently, for each branch, we utilize CNN backbones to extract dynamic spatial depth features from video frames and log-spectrograms, followed by the use of Temporal Convolutional Networks[1] to further learn spatiotemporal encoding. Lastly, these features are concatenated and fed into a Transformer structure to capture temporal information for downstream tasks. During training, we employ a strategy of large-window resampling, where the window size determines the amount of contextual information the model considers for predicting each time step.
The main contribution of the proposed method can be summarized as:
1. We introduce a novel approach for fusing multimodal data, leveraging pre-trained audio and video backbones to extract dynamic features from both visual and audio modalities.
2. Utilizing Temporal Convolutional Networks (TCN), we further learn spatiotemporal encoding to capture temporal dependencies and patterns in the data.
3. We utilize a Transformer structure to effectively capture temporal information for downstream tasks, ensuring robust performance in real-world scenarios.
4. During training, we employ a large-window resampling strategy to determine the amount of contextual information considered for predicting each time step, enhancing the model’s ability to understand long-term dependencies in the data.
2 Related Work
In the field of affective behavior analysis, multimodal feature fusion is a key issue. Most studies[20, 25] adopt multimodal methods to improve model performance by fusing visual and audio features. In previous research, researchers have proposed various fusion methods, such as directly fusing multimodal features into a common feature vector for analysis[19], or separately analyzing the features of each modality and then fusing the final outputs to obtain better prediction results[32], or even combining the above two fusion methods to achieve better results[17, 15]combined visual and audio information in videos, constructing a dual-stream network for emotion recognition, achieving high performance. Currently, convolution-based methods[22, 34] have made significant progress. At the same time, the application of Transformer in multimodal learning has become mainstream, [21, 7, 11]achieving excellent performance using the Transformer structure. In terms of audio features[33, 24, 16], energy features, temporal features, etc., are widely used to improve the ability of emotion recognition.
Recently, researchers have proposed many new methods and frameworks[4, 27, 28, 31], including tasks such as AU detection, expression recognition, and VA estimation. Some of these methods leverage the correlation between VA and AU or VA and EXPR, proposing multi-task frameworks. These methods can extract complementary information from other tasks. To address the high cost of annotating emotion/AU/VA labels from real-world facial images, some researchers have proposed using self-supervised learning (SSL) methods to leverage knowledge from existing large-scale unlabeled data. These methods provide new avenues for improving the accuracy of affective analysis.
In conclusion, multimodal features play an important role in affective behavior analysis, and various fusion methods and self-supervised learning methods provide new ideas and approaches for improving the accuracy and efficiency of emotion recognition.
Figure 1: Our proposed framework for VA estima-tion.
3 Proposed approach
In this section, we focus on introducing our method. To better accomplish the VA estimation task, we have designed an efficient network. As shown in Figure 1, we use a CNN backbone to extract dynamic spatial depth features from video frames and audio. Subsequently, we utilize TCN to further learn spatiotemporal encoding. Afterward, we concatenate the video and audio features and feed the output into a fully connected layer to apply to the VA estimation task.
3.1 Feature Extraction
VGGish:VGGish[9] is a pre-trained deep neural network model developed by Google, specifically designed for audio feature extraction. It is commonly used for tasks such as audio classification, audio event detection, and other audio-related tasks. The model architecture is based on the VGG network architecture, which consists of multiple convolutional layers followed by max-pooling layers. Its output is a 128-dimensional feature vector that can be used for speech-related tasks.
MFCC:MFCC[3] is a widely used method for audio feature extraction in tasks such as speech recognition and audio processing. It was proposed by American engineer Harvey Fletcher and Japanese engineer Takeo Yamazaki in the 1960s. Based on the perceptual characteristics of the human ear to sound frequencies, MFCC transforms audio signals into a set of feature vectors to represent the spectral characteristics of the audio. The process of extracting MFCC involves segmenting the audio signal into frames, applying windowing, performing a fast Fourier transform to compute the power spectrum, applying a Mel filterbank, taking the logarithm, and finally applying discrete cosine transform to obtain the MFCC coefficients. These coefficients can be used to train machine learning models, such as speech recognition and speaker recognition, for the analysis and processing of audio data. Its output is a 39-dimensional feature vector that can be used for speech-related tasks.
IResNet-50:IResNet-50[2] is a variant of the ResNet[8]architecture, which was introduced by Microsoft Research in their paper ”Deep Residual Learning for Image Recognition” by Kaiming He et al. IResNet-50 incorporates the idea of residual connections to enable the training of very deep neural networks more effectively. In this study, we trained IResNet-50 using the AffectNet[18] dataset and obtained a 512-dimensional visual feature vector.
Table 1: The dimensions of features.
Feature Modality Dimension Description
VGGish 128 A
MFCC 39 A
IResNet-50 512 V
3.2 Preprocessing
Our preprocessing approach is primarily based on the solution proposed by Su Zhang, whoes were last year’s third-place winners[30]. In our experiments, the input consists of one visual feature and two audio features, but their sizes are typically different, and there may even be significant discrepancies. The sizes of the features are shown in Table 1. The visual preprocessing first resizes all images to a size of 48×48×3, based on the cropped-aligned image data provided by the organizer. For each trial in the training or validation set, its length is determined by the number of rows in the annotation text file, excluding those marked as -5. In the test set, the length is determined by the frame count of the original video. A zero matrix of size N×48×48×3 is initialized, and then each row is assigned to the corresponding jpg image if it exists. The audio preprocessing converts all videos to mono with a 16K sampling rate in WAV format initially. Then, the logmelspectrogram is extracted using preprocessing code from the Vggish repository, with the hop length specified to synchronize with other modalities and annotations. Emotion label processing involves excluding rows containing -5. To ensure consistency in length between features and annotations, the feature matrix is either padded repeatedly (using the last feature points) or trimmed (from the rear), depending on whether the feature length is shorter or longer than the trial length.
3.3 Training
In this study, we employed a six-fold cross-validation method to fully utilize the AffWild2 database. The database consists of 360 training trials, 72 validation trials, and 162 testing trials, covering various emotional expressions and contexts. Through cross-validation, we were able to leverage the information in the dataset effectively, reducing the model’s dependence on specific data distributions, thereby better assessing the performance and robustness of the model. Given the video frames $V_{i}$ and the corresponding audio segments $A_{i}$ obtained through preprocessing, we employ pre-trained video feature extraction models (IResNet-50) and audio feature extraction models (VGGish) to extract visual features $F_{\text{vis}}^{i}$ and audio features $F_{\text{aud}}^{i}$ for each frame separately. Subsequently, we send them to TCN modules for processing to capture the temporal and spatial correlations between them. As shown in Equation 1, we concatenate the extracted visual frame feature branch and the audio feature branch. Next, the processed features are fed into a Transformer encoder structure, which helps learn long-range dependencies between features and improves the performance and generalization ability of the model. The Transformer encoder consists of four encoder layers, each with a dropout rate of 0.4. The output of the encoder is then passed through a fully connected layer, a Batch Normalization layer, and another fully connected layer. Finally, a tanh activation function is used for prediction, yielding the Valence (V) or Arousal (A) values for each video frame. This design enhances the model’s expressive power and stability, enabling it to better adapt to various tasks and data distributions. When computing the loss, we utilize the CCCLoss as shown in Equation 2.Due to the frame-wise prediction process, we adopt a smoothing strategy to enhance the stability of the model. We observed occasional omissions of frames in the cropped face images extracted from the videos. Since video frames are continuous, we replace the missing frames with the last available frame.
$f=\text{concat}(\hat{f}_{1},\hat{f}_{2},\hat{f}_{3})$ (1)
$\text{CCCLoss}=1-CCC(\hat{v}_{\text{frame}_{i}},v_{\text{frame}_{i}})+1-CCC(\hat{a}_{\text{frame}_{i}},a_{\text{frame}_{i}})$ (2)
4 Experiment
In this section, we provided an overview of the dataset, evaluation protocol, and experimental results used in the study.
4.1 Dataset and Evaluation
Affwild2 dataset. Aff-Wild2 is an extension of the Aff-Wild dataset for affect recognition. It approximately doubles the number of included video frames and the number of subjects; thus, improving the variability of the included behaviors and of the involved persons. Aff-Wild2 is a significant research asset, being the only database annotated for all three main behavioral tasks in the wild. It is a large-scale database and the first one to feature AU annotations alongside both audio and video. Annotations in Aff-Wild2 are frame-based, encompassing seven basic expressions, twelve action units, as well as valence and arousal. In total, Aff-Wild2 comprises 564 videos, around 2.8 million frames, involving 554 subjects. These subjects represent diverse demographics in terms of age, ethnicity, and nationality, while also presenting a wide array of environmental and situational variations.
Evaluation.To estimate the VA, we computed the Concordance Correlation Coefficient (CCC) separately for arousal and valence. The evaluation metric for this competition is represented by Equation 3.
$P=0.5\times(CCC_{\text{arousal}}+CCC_{\text{valence}})$ (3)
4.2 Results on Validation set
For the estimation of VA, we evaluate the models based on the average CCC values for valence and arousal. To enhance the model’s generalization, we further conduct a six-fold cross-validation on randomly segmented annotated data. Detailed experimental results are provided in Table 2. This table presents the Valence and Arousal scores for different validation sets (fold-0 to fold-5). Each validation set corresponds to a specific cross-validation fold. Each row represents a validation set, including its Valence and Arousal scores. The last row shows the best Valence and Arousal scores obtained across all folds, highlighted in bold.
Table 2: Valence and Arousal Scores for Different Validation Sets
Val Set Valence Arousal
fold-0 0.5121 0.5507
fold-1 0.5789 0.5992
fold-2 0.5547 0.5862
fold-3 0.5743 0.5959
fold-4 0.5971 0.6156
fold-5 0.6123 0.6589
5 Conclusion
This paper introduces our submission for the VA estimation task in ABAW6 competition. We conducted preprocessing on video frames and audio segments to extract visual and audio features, constructing a comprehensive model for VA (Valence-Arousal) estimation task. Utilizing TCN modules, we captured the temporal and spatial correlations between features. Finally, we fed the extracted features into a Transformer encoder structure to learn long-range dependencies and enhance the model’s performance and generalization ability.